OpenStudy (anonymous):

please give me some examples for the closure property in this set: The odd numbers is closed under multiplication.

OpenStudy (anonymous):

take two general odd numbers , 2k+1 and 2n+1 (2k+1)(2n+1)

OpenStudy (anonymous):

= 4kn + 2 (k+n ) + 1 = 4kn +2k +2n + 1 = 2 [ kn + k +n ] + 1 let kn +k +n = C ( another constant )

OpenStudy (anonymous):

therefore (2k+1)(2n+1) = 2C +1 ( the general form of an odd number ) there

OpenStudy (anonymous):

not an example, but a proof, good enough

OpenStudy (anonymous):

made a mistake above should be : 2 [ 2kn + k +n ] + 1

OpenStudy (anonymous):

but the argument is the exact same, that expression in the bracket is a sum of constants, and can therefore be represented by a single constant.